[figure: 9.7 cm]

Magnetostriction in superconducting MgB_{2}

A. Nabialek^{a}, B. Kundys^{a}, Yu. Bukhantsev^{a}, S. Vasiliev^{a}, A. Wiśniewski^{a},

J. Jun^{b}, S.M. Kazakov^{b}, J. Karpiński^{b}, A. Szewczyk^{a}, M. Gutowska^{a}, K. Piotrowski^{a}, H. Szymczak^{a} a Institute of Physics of the Polish Academy of Sciences, Al. Lotników 32/46, 02-668, Warsaw, Poland b Solid State Physics Laboratory, ETH, 8093 Zürich, Switzerland

The magnetostriction measurements have been performed at various temperatures for superconducting MgB_{2} using three terminal capacitance technique in magnetic field up to 12 T. The obtained results were analyzed in frames of the pinning induced model of magnetostriction. The performed scaling analysis of obtained experimental data suggests that the pinning mechanism responsible for magnetostriction in MgB_{2} is unique over broad temperature region below the critical temperature.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Yuriy Bukhantsev

Address for correspondence : Institute of Physics Polish Academy of Sciences, Al. Lotnikow 32/46, 02-668, Warsaw, Poland

Email address : bukha@ifpan.edu.pl